### v1

# Warehouse storage and retrieval optimization via clustering, dynamic systems modeling, and GPU-accelerated routing

Abstract.
This paper introduces a warehouse optimization procedure aimed at enhancing the efficiency of product storage and retrieval. By representing product locations and order flows within a time-evolving graph structure, we employ unsupervised clustering to define and refine compact order regions, effectively reducing picking distances. We describe the procedure using a dynamic mathematical model formulated using tools from random dynamical systems theory, enabling a principled analysis of the system’s behavior over time even under random operational variations. For routing within this framework, we implement a parallelized Bellman-Ford algorithm, utilizing GPU acceleration to evaluate path segments efficiently. To address scalability challenges inherent in large routing graphs, we introduce a segmentation strategy that preserves performance while maintaining tractable memory requirements. Our results demonstrate significant improvements in both operational efficiency and computational feasibility for large-scale warehouse environments.
1. Introduction
The exponential growth of the e-commerce sector has necessitated significant advancements in warehouse management strategies. Increasing consumer demand for expedited order fulfillment has placed pressure on logistics operations, and major retailers like Amazon, H&M, and Walmart have made substantial investments in automation and AI-driven warehouse management systems (WMS) to enhance efficiency. One of the primary cost drivers in warehouse operations is order picking, accounting for approximately 60% of total warehouse expenses [14]. It directly impacts labor costs, order fulfillment speed, and overall supply chain performance. A well-optimized picking strategy minimizes travel distance while maintaining flexibility to accommodate variations in order composition over time. Since order-picking routes can be represented as graphs, where nodes correspond to product locations, determining an optimal route for each order is a combinatorial optimization problem. To further enhance efficiency, order clustering strategies based on product location and demand frequency have been proposed, allowing frequently co-ordered products to be relocated closer together, thereby reducing long-term travel distances [15].
Shortest path algorithms play a key role in optimizing order picking. While Dijkstra’s algorithm [7] is a well-known approach, its sequential nature limits its efficiency for large-scale problems. The Bellman-Ford algorithm [3, 8], in contrast, is better suited for parallel computing [10] and has been successfully implemented on GPUs to accelerate shortest path computations [5]. However, in real-world warehouse settings, order picking involves combinatorial optimization, where factorial growth in possible routes makes exhaustive evaluation impractical. In fact, finding a shortest path for picking operations is a version of the traveling salesman problem (TSP). Since the TSP is NP-complete [6], approximate methods are often necessary in practice.
In addition, warehouse managers frequently encounter unpredictable demand patterns and inventory fluctuations, challenges that static optimization methods struggle to address effectively. Instead, stochastic modeling approaches have been shown to significantly enhance logistics decision-making, particularly regarding inventory positioning and order fulfillment [9]. This is exemplified in Amazon’s fulfillment centers where stochastic modeling is incorporated in their robotic picking algorithm to balance pick efficiency and order cycle times. While clustering in the traditional data science sense isn’t explicitly used, the algorithm conceptually groups orders based on various criteria to optimize efficiency [1].
Derived from the analysis above, this study proposes an advanced warehouse optimization framework that integrates graph databases, unsupervised clustering, and GPU-accelerated route optimization with the Bellman-Ford algorithm. Unlike conventional WMS that rely on static zone allocations, our approach continuously repositions products based on demand fluctuations. This research extends prior work in warehouse route optimization and clustering methodologies [16, 4] by employing a random dynamical systems (RDS) framework that provides a rigorous mathematical foundation for analyzing storage and retrieval efficiencies and modeling uncertainty and variability. By leveraging GPU-based computing, our system can handle large-scale warehouses and high order volumes more effectively.
The system employs clustering on two levels: (1) Order clustering, which reorganizes storage layouts to concentrate frequently co-ordered items and minimize long-term travel distance, and (2) Picking node clustering, which reduces computational complexity in route optimization while maintaining near-optimal results. These techniques are analyzed mathematically in Sections 2 and 3 and evaluated through numerical experiments in Section 4. By continuously adapting storage configurations and applying combinatorial optimization strategies, this approach balances short-term efficiency gains with long-term structural improvements, ensuring that warehouse operations remain effective even amid fluctuating order patterns.
To describe the implementation of our system in full detail we develop a mathematical model found in Section 2. The approach can be summarized as follows:
Products in the warehouse are organized using a graph database, which allows for real-time tracking of their positions. This system ensures that the location of each product is always known, facilitating efficient management of stock levels and movement. Incoming orders are clustered based on their similarity and proximity, grouping orders that share common product types or are located near each other within the warehouse. This clustering is also taken into account for calculating picking routes, which minimizes the travel distance for picking multiple orders, thereby increasing efficiency.
For each cluster of orders, a cluster center is determined. This center represents the optimal location within the warehouse for storing products that are frequently included in the cluster’s orders. The calculation of the cluster center takes into account the positions of all products in the cluster and aims to minimize the overall distance required for picking. When a product position is emptied during the picking process, the system triggers a reorganization of the warehouse. The product is restocked at a new position that is closer to the relevant cluster center, ensuring that products are always stored in positions that reduce the travel distance for future picking operations.
2. Mathematical model
We describe our warehouse optimization method and our numerical experiments using dynamical systems theory. We first introduce the state space (describing the different states of the warehouse) and give the definition of orders.
2.1. The state space
We consider a situation where $N$ types of articles $a_{1},a_{2},\ldots,a_{N}$ are stored in a warehouse at nodes arranged in three-dimensional space representing racks with shelves at different heights. Each shelf on a rack constitutes its own node, and each node is identified with an element in a three-dimensional array having integer coordinates
$(i,j,k)\in[1,n_{x}]\times[1,n_{y}]\times[1,n_{z}]\subset\mathbb{N}^{3},$ (1)
where =$\mathbb{N}=\{0,1,2,\ldots\}$ is the set of natural number, and $n_{x}\times n_{y}\times n_{z}$ is the total number of nodes in the warehouse. We assume there is at most one article type $a_{n}$ at each node (some nodes being empty), and identify each article type $a_{n}$ with the corresponding ordinal $n\in[1,N]$. Empty nodes are assigned article type 00.
For each article type $n\in[1,N]$, let $M_{n}$ be the maximum number of parcels that can be stored at an individual node. We assume this number is independent of which node is being considered; it only depends on what type of article is stored at that node. Introduce the space $\mathcal{A}=[0,N]^{n_{x}\times n_{y}\times n_{z}}$ of three-dimensional arrays. Each array $A=(a_{ijk})_{ijk}\in\mathcal{A}$ corresponds to one particular choice of how to distribute article types in the warehouse. (The array with only zeros represents an empty warehouse with no article types, and the array with only ones is a warehouse that only contains articles of type $a_{1}$.) Each array element $a_{ijk}$ is thus an integer $n$ between 00 and $N$, representing the fact that the node at $(i,j,k)$ contains article type $a_{n}$.
In other words, for each fixed $A=(a_{ijk})_{ijk}\in\mathcal{A}$, the node $(i,j,k)$ contains articles of type $a_{ijk}\in[0,N]$. As described above, the maximum number of such articles at one node is $M_{a_{ijk}}$. For each $A=(a_{ijk})_{ijk}\in\mathcal{A}$ we now consider the fiber over $A$,
$F_{A}\mathcal{A}=\{M:M=(m_{ijk})_{ijk},\ m_{ijk}\in[0,M_{a_{ijk}}]\}.$
This describes all the possible inventory balances $M$ (number of parcels of each article type) for a given article type distribution $A$. The state space is then the fibration defined by the (disjoint) union
$X=\bigcup_{A\in\mathcal{A}}\{A\}\times F_{A}\mathcal{A}=\{(A,M):A\in\mathcal{A},\ M\in F_{A}\mathcal{A}\},$
where $A$ varies over all the possible article type distributions. States in $X$ will be denoted by $x$, so that $x=(A,M)\in X$. We may of course also represent $X$ by attaching to each article type the corresponding inventory balance, so that
$X\cong\{(a_{ijk},m_{ijk})_{ijk}:a_{ijk}\in\mathcal{[}0,N],\ m_{ijk}\in[0,M_{a_{ijk}}]\},$
and we will not distinguish between these two representations.
2.2. Orders
An order $o$ is a vector
$o=(m_{1},\ldots,m_{N})\in\mathbb{N}^{N}$
where $m_{n}\in\mathbb{N}$ is the number of parcels of article $a_{n}$ contained in the order. (Here we shall by an order mean any order – individual customer orders or a collection of customer orders – that has been decided should be picked from the warehouse.) Given a state $x=(A,M)$ and an order $o$, the Order Processing algorithm (defined in Algorithm 1 in Appendix A below) and a route optimization algorithm (which involves the proposed clustering step based on similarity and proximity, see §2.4) provide the nodes from which the order should be collected, along with the picking route between these nodes. The first part is described by a map $w$ given by
$w:(x,o)\mapsto w(x,o)=(w_{ijk})_{ijk}\in\mathbb{N}^{n_{x}\times n_{y}\times n_{z}},$ (2)
where $w_{ijk}$ is the number of parcels that should be collected from node $(i,j,k)$, and the total number of parcels in the order is $\Sigma_{i,j,k}w_{ijk}=m_{1}+\ldots+m_{N}$. The Update Stock algorithm (Algorithm 2) then calculates the intermediate state
$(A,M^{\prime})=(A,M-w(x,o)).$ (3)
The Check Stock algorithm (Algorithm 3) verifies if there is sufficient stock to fill the order without picking the last parcel on a shelf, i.e., whether $m_{ijk}>w_{ijk}$ for all $(i,j,k)$. If this is not the case, i.e., if $m_{ijk}\leq w_{ijk}$ for some $(i,j,k)$ such that $w_{ijk}\geq 1$, then the Check Stock and Move Articles algorithm of the warehouse optimization routine (Algorithm 4) computes a new warehouse location $(i^{\prime},j^{\prime},k^{\prime})$ at which to restock article $a_{ijk}$ (collecting from the new location after emptying the balance at the old location), and the article array $A$ is updated to $A^{\prime}$ in which $a_{ijk}$ is replaced by 00, and $a_{i^{\prime}j^{\prime}k^{\prime}}$ is changed from 00 to $a_{ijk}$. This gives an updated state
$x^{\prime}=f(x,o):=(A^{\prime},M^{\prime}),$ (4)
which is the new state of the warehouse after having completed picking of the order $o$. If no nodes become empty when filling the order then the intermediate state (3) is taken as the the new state $x^{\prime}$ in (4).
2.3. Random dynamical systems
We now recall the notion of a random dynamical system (RDS), a symbiosis of ergodic theory and topological dynamical systems. An advantage of the RDS framework is that it also applies in situations where orders have a random element. Standard references are Arnold [2] and Strogatz [17].
We use discrete time $n\in\mathbb{N}$, where =$\mathbb{N}=\{0,1,2,\ldots\}$ as before is the set of natural numbers. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space. We take $\Omega$ to be the shift space consisting of sequences $\omega=(o_{1},o_{2},\ldots)$ of orders $o_{n}$, and let $\theta:\Omega\to\Omega$ to be the shift operator $\theta\omega=(o_{2},o_{3},\ldots)$. This models a sequence of orders $\omega\in\Omega$, placed with probability $\mathbb{P}$, coming in to the warehouse to be filled one after the other, with $\theta$ driving the flow of orders. Let $X$ be the state space describing different states of the warehouse. Let $F:\omega\mapsto F_{\omega}$ be a random variable from $\Omega$ to the set of maps on $X$, i.e., for all $\omega\in\Omega$, $F_{\omega}$ is a map $F_{\omega}:X\to X$. Consider the skew product defined by $\Theta(\omega,x)=(\theta\omega,F_{\omega}(x))$. Iterating gives
$\Theta^{n}(\omega,x)=(\theta^{n}\omega,F_{\theta^{n-1}\omega}\circ F_{\theta^{n-2}\omega}\circ\cdots\circ F_{\omega}(x))$
for $n\geq 1$, where $\circ$ denotes composition of functions. The second coordinate is a random dynamical system over $\theta$.
We let $f(x,o)$ be the function in (4) describing how the state $x$ evolves to $x^{\prime}=f(x,o)$ by filling the order $o$ once, then, if necessary, restocking articles at new locations as determined by the optimization routine. With $\omega=(o_{1},o_{2},\ldots)$ we then let $F_{\omega}(x)=f(x,o_{1})$, so that $F_{\theta^{n-1}\omega}(x)=f(x,o_{n})$ for $n\geq 1$. The simplest example is when $\omega=(o,o,\ldots)$, corresponding to the same order $o$ being placed over and over (see Experiment 1 in §4 below). To account for small differences in individual orders we will also consider the random example when $\omega=(o_{1},o_{2},\ldots)$ and the orders $o_{n}$ are independent, identically distributed perturbations of some fixed order $o$ (see Experiments 2 and 3 in §4). In this case $\Theta^{n}(\omega,x_{0})$ describes how the (initial) warehouse state $x_{0}$ evolves to
$x_{n}=F_{\theta^{n-1}\omega}\circ\cdots\circ F_{\omega}(x_{0})=f(\bullet,o_{n})\circ\cdots\circ f(\bullet,o_{1})(x_{0})$ (5)
by filling the first $n$ orders in $\omega$, having potentially restocked articles after each order at new locations as determined by the optimization routine. $\Theta^{n}(\omega,x_{0})$ also contains the information that the next order to be filled is the first entry in $\theta^{n}\omega=(o_{n+1},o_{n+2},\ldots)$, that is, $o_{n+1}$. We also note that in this case, the random difference equation
$x_{n}=F_{\theta^{n-1}\omega}(x_{n-1}),\quad n\geq 1,\quad x_{0}\in X,$
generates a Markov chain $(x_{n})_{n}$ with transitional probability $P(x,B)=\mathbb{P}\{\omega:F_{\omega}(x)\in B\}$, cf. [13]. Iteration of the picking and restocking process for a sequence of orders $o_{1},o_{2},\ldots$ in this way is handled by the Main Order Processing algorithm (Algorithm 5).
2.4. Routes and clusters
Given a state $x$ and an order $o$, a route is an ordering of the elements in the set of picking nodes
$S_{\rm pick}=S_{\rm pick}(x,o)=\{(i,j,k):w_{ijk}\neq 0\}$
determined by (2). Routes are determined so that picking is completed from nodes on the same rack first (i.e., nodes that have the same $(i,j)$-coordinates) before moving on to the next rack. The stops along the route can then be described as a given ordering of the elements in the set of stopping nodes
$S_{\rm stop}=S_{\rm stop}(x,o)=\{(i,j):\text{ there exists $k$ such that }(i,j,k)\in S_{\rm pick}\},$
and the number of stops is
$n=\#S_{\rm stop}.$ (6)
At stopping nodes with more than one picking node, we can pre-determine the order in which picking is carried out – e.g., top to bottom – so this does not have to not included in the route optimization. We will therefore not distinguish between picking routes and routes between stopping nodes.
Figure 1. Each of the top panels show one of three clusters resulting from $K$-means clustering applied to an order with 200 article types given a randomized initial warehouse state $x_{0}$. The nodes in the warehouse have coordinates $(i,j,k)$ with $i,j,k=1,\ldots,10$. Stopping nodes are indicated with color coding describing how many picking nodes a given stopping node contains. Each of the bottom panels show one of three clusters resulting from $K$-means clustering applied to the same order being picked for the 100:th time, that is, given the warehouse state $x_{99}$ as in (5), obtained after 99 picking iterations. In the top panels, all three clusters are spread out over the entire warehouse with much overlap between clusters. In contrast, the clusters in the bottom panels are concentrated and much better separated.
As explained in the introduction, we propose to group orders into clusters based on similarity and proximity to reduce computational complexity while still optimizing picking. The separation and concentration of the clusters is improved as time progresses by taking clustering into account when restocking. The algorithm restocks an article at an empty picking node whose $(x,y)$-coordinate (i.e., stopping node) is closest to the cluster center of the corresponding cluster. This reduces the area over which each cluster is spread out. Figure 1 shows an example of this phenomenon, which we will investigate further in Section 4 by performing several numerical experiments. Each of the top panels in Figure 1 show one of three clusters resulting from applying $K$-means clustering on a picking order $o$ consisting of 200 article types, given a randomized initial warehouse state $x_{0}$. The nodes in the warehouse have coordinates $(i,j,k)$ with $i,j,k=1,\ldots,10$. The stopping nodes in each cluster are indicated with color coding describing how many picking nodes there are at each stopping node. The three clusters are all more or less spread out over the entire square. In contrast, the bottom panels show $K$-means clustering when the same order $o$ is going to be picked for the 100:th time; the warehouse state is then $x_{99}$ with $x_{n}$ as in (5), obtained after applying the resorting algorithm having already picked this order 99 times. Article types have now been restocked at new positions that are closer to the relevant cluster center, and clusters are clearly distinguishable and well separated.
To make this mathematically precise we introduce the following definition. Let $m$ be the number of clusters. Given a state $x$ and an order $o$, a cluster is a subset of $S_{\rm pick}$ which we will denote by
$S_{\rm pick}(\ell)=S_{\rm pick}(\ell;x,o)\subset S_{\rm pick},\quad 1\leq\ell\leq m,$
such that the $S_{\rm pick}(\ell)$ are mutually disjoint for $1\leq\ell\leq m$ and
$S_{\rm pick}=\bigcup_{\ell=1}^{m}S_{\rm pick}(\ell).$
The stops in cluster $\ell$ are
$S_{\rm stop}(\ell)=S_{\rm stop}(\ell;x,o)=\{(i,j):\text{ there exists $k$ suchthat }(i,j,k)\in S_{\rm pick}(\ell)\}.$ (7)
Note that the sets $S_{\rm stop}(1),\ldots,S_{\rm stop}(m)$ are not mutually disjoint in general, which means that a rack in the warehouse can belong to more than one cluster. In (7), the number of $k$’s such that $(i,j,k)\in S_{\rm pick}(\ell)$ is the picking frequency at stop $(i,j)$ for cluster $\ell$. By taking picking frequency into account, clusters can be represented in two dimensions as in Figure 1, where the frequency is represented by the color coding of each stopping node.
Remark.
Note that we define clusters mathematically as a subset of the picking nodes for a given order, as opposed to dividing the picking nodes from a collection of orders into different groups or clusters. But since an order here can refer to the total collection of individual customer orders to be picked at a certain time our definition covers also the second scenario.
Because we will study routes between stopping nodes, we define cluster centers as the expected value of the location of stopping nodes in each cluster, taking picking frequency into account. The cluster centers are then given for cluster $\ell$ by $(\bar{x}(\ell),\bar{y}(\ell))$ where
$\bar{x}(\ell)=\frac{\sum_{(i,j,k)\in S_{\rm pick}(\ell)}i}{\#S_{\rm pick}(\ell)},\quad\bar{y}(\ell)=\frac{\sum_{(i,j,k)\in S_{\rm pick}(\ell)}j}{\#S_{\rm pick}(\ell)},$
with covariance matrix
$\Sigma(\ell)=\begin{pmatrix}\sigma_{xx}^{2}(\ell)&\sigma_{xy}^{2}(\ell)\\ \sigma_{xy}^{2}(\ell)&\sigma_{yy}^{2}(\ell)\end{pmatrix}.$ (8)
Here the variances and covariance are given by
$\sigma_{xx}^{2}(\ell)=\frac{\sum_{(i,j,k)\in S_{\rm pick}(\ell)}(i-\bar{x}(\ell))^{2}}{\#S_{\rm pick}(\ell)},$
$\sigma_{yy}^{2}(\ell)=\frac{\sum_{(i,j,k)\in S_{\rm pick}(\ell)}(j-\bar{y}(\ell))^{2}}{\#S_{\rm pick}(\ell)},$
$\sigma_{xy}^{2}(\ell)=\frac{\sum_{(i,j,k)\in S_{\rm pick}(\ell)}(i-\bar{x}(\ell))(j-\bar{y}(\ell))}{\#S_{\rm pick}(\ell)}.$
To study how clusters evolve as we apply our resorting procedure to a sequence of orders $o_{1},o_{2},\ldots$ we let $\Sigma_{n}(\ell)$ be the covariance matrix of cluster $\ell$ as in (8) defined for a state $x_{n-1}$ and order $o_{n}$, $n\geq 1$. Here $x_{n}$ is given by (5), so that $x_{n}$ is the warehouse state obtained after filling the first $n$ orders in the sequence $o_{1},o_{2},\ldots$. This will be used in Section 4 where we illustrate the time evolution of the mean and variance of clusters by plotting bivariate Gaussian probability density functions with corresponding mean and covariance matrices as $n$ increases, see Figures 3, 4 and 5.
3. Route Optimization
In addition to clustering orders for reorganizing storage layouts, this paper presents a procedure for optimizing order picking routes by using clustering also within a given picking order. In the language of our mathematical model, we recall that a picking route is a route with total number of stops in the $xy$-plane given by $n$ in (6). If there are more than one picking node at a picking stop (cf. §2.4) we pre-define the picking order for such nodes so this does not have to be included when calculating the route. However, we recall that the frequency of picking at each stop does matter, and that the restocking routine biases toward increased frequency. Over time (i.e., upon iteration of the resorting algorithm) this leads clusters to become more concentrated with smaller area and fewer stops, which in turn leads to a reduction in both travel distance and computational complexity. However, finding an optimal route can be a daunting task even after the number of stops have been reduced in this way, so in this section we discuss a technique to reduce computational complexity while maintaining near optimal results when calculating the route at each iteration of the resorting algorithm. We showcase the technique at the end of Section 4.
For large orders, where the computational complexity of route optimization can be $O(n!)$, we propose to use a GPU-based implementation of the Bellman-Ford algorithm to compute the route. This approach allows for efficient handling of large datasets by significantly reducing computational time. For clarity of exposition we will assume that the entrance and exit are the same in the warehouse, and that these are not part of the route. Routes will then not be closed curves, and we will treat two routes as identical if they differ only in their travel direction. The number of possible routes between $n$ picking stops (i.e., stopping nodes) is then $\frac{1}{2}n!$. To reduce computing costs we propose to use unsupervised machine learning to divide the picking stops into $m$ clusters, then finding optimal routes in each cluster, and finally patching these routes together. This has the possibility of substantially reducing the computational complexity. For simplicity we begin with a naive discussion that doesn’t leverage any other sophisticated algorithms to reduce complexity.
Theorem.
Let $n$ be the number of picking stops, and let these be divided into $m$ pairwise disjoint clusters, each containing $n_{1},\ldots,n_{m}$ picking stops, where $n_{1}+\ldots+n_{m}=n$. If an undirected non-closed route has been determined for each cluster then the total number of possible global undirected routes connecting them is $m!\times 2^{m-1}$. The total cost of also finding optimal routes in each individual cluster is therefore
$m!\times 2^{m-1}+\frac{1}{2}(n_{1}!+\ldots+n_{m}!).$
Proof.
Suppose first that an undirected route with two boundary nodes has been found for each cluster. We determine the computational complexity of finding the best global route going through all clusters along these routes using a combinatorial argument. First, a starting cluster is chosen which can be done in $m$ ways. From that cluster, a global starting point should be determined from the pair of that cluster’s boundary nodes, which can be done in $2$ ways. Next, a second cluster is chosen which can be done in $m-1$ ways. From that cluster, a starting point connecting with the first cluster’s end point should be determined from the second cluster’s pair of boundary nodes, which can be done in $2$ ways. Continuing this way we see that there are $m!\times 2^{m}$ ways to choose a global route with a start and end. Hence, there are half as many global undirected routes, namely $m!\times 2^{m-1}$.
Now, for a cluster with $n_{j}$ picking stops there are $\frac{1}{2}n_{j}!$ ways to choose an undirected route. This is independent of the other clusters. Adding these to the cost of choosing a global route completes the proof. ∎
Figure 2. An order with 15 stopping positions clustered into 3 regions (left), and the reduced routes for each cluster represented by two boundary nodes (right).
The reduction in computational cost described by the theorem is largest when $n_{j}\sim(n/m)$ for $j=1,\ldots,m$. To give the reader an idea of the magnitude of the reduction in this case, consider a modest $n=12$. Then $12!\approx 5\times 10^{9}$, while if $n_{j}=n/m$ for $j=1,\ldots,m$ then
$m!\times 2^{m-1}+\frac{1}{2}(n_{1}!+\ldots+n_{m}!)=\begin{cases}724&m=2,\\ 60&m=3,\\ 204&m=4.\end{cases}$
As can be seen, this approach significantly reduces the computational complexity. Although it does not guarantee an optimal route, it ensures that a route close to the optimal one can be achieved within a reasonable computational time.
The case when $n=15$ and $m=3$ is shown in Figure 2. Then $15!\approx 10^{12}$, while $n_{1}=n_{2}=n_{3}=5$ and
$m!\times 2^{m-1}+\frac{1}{2}(n_{1}!+\ldots+n_{m}!)=204.$
Of course, we have been comparing the results of the theorem above to a brute force approach for finding optimal routes without employing sophisticated algorithms to reduce computational complexity. However, using for example the Held & Karp algorithm [11] reduces the time complexity from $\mathcal{O}(n!)$ to $\mathcal{O}(n^{2}\cdot 2^{n})$, which, with $n=12$, is still 3 orders of magnitude higher than the examples above – and this without using a corresponding algorithm to find optimal routes within clusters. Hence, the gains in time complexity from the approximate method are still substantial.
4. Numerical experiments
Here we perform numerical experiments to evaluate the performance of the resorting algorithm for a warehouse model described in Section 2 with the following parameter settings:
  1. (1)
    The warehouse contains $10\times 10\times 10$ picking nodes located at $(i,j,k)\in[1,10]\times[1,10]\times[1,10]$. This means that $n_{x}=n_{y}=n_{z}=10$ in (1).
  2. (2)
    The number of article types is $N=890$, and we assume that the maximum number of articles of each article type that a node can stock is $10$, so $M_{a_{ijk}}=10$ for all $(i,j,k)$. Thus, the state space $X$ is
    $X=\{(a_{ijk},m_{ijk})_{ijk}:a_{ijk}\in[0,890],\ m_{ijk}\in[0,10]\}.$
  3. (3)
    The initial warehouse state $x_{0}$ is randomized with the following constraints: It contains $11$ $xy$-coordinate pairs $(i_{0},j_{0}),(i_{1},j_{1}),\ldots,(i_{10},j_{10})$ such that the nodes $(i_{0},j_{0},k),\ldots(i_{10},j_{10},k)$ are empty for all $k\in[1,10]$. This corresponds to $11$ empty racks, for a total of $110$ empty nodes. A unique article type $a_{ijk}\in[1,890]$ is assigned to each of the remaining $890$ nodes, and we assume that the initial warehouse state is fully stocked, so the balance at each nonempty node $10$.
  4. (4)
    An order $o$ to be picked is a collection of $20$ individual customer or purchase orders. A purchase order consists of $10$ article types with between $1$ and $10$ parcels of each type. The $20$ purchase orders are mutually disjoint, so $o$ consists of $200$ article types. In each picking iteration, the $20$ purchase orders are divided into three groups using $K$-means clustering based on where the articles are stored in the warehouse at that time. Each such group is a cluster. An example can be seen in Figure 1.
4.1. Experiment 1
We first perform picking and resorting when the same order is placed over and over, that is, we let $\omega=(o,o,\ldots)$ and compute the warehouse state
$x_{n}=F_{\theta^{n-1}\omega}\circ\cdots\circ F_{\omega}(x_{0})=f(o,\bullet)\circ\cdots\circ f(o,\bullet)(x_{0}),\quad n\geq 1,$
obtained after filling the order $n$ times, having restocked articles at potentially new locations as determined by the optimization routine. This is depicted in Figure 1, and the experiment shows that continued application of the picking and resorting procedure leads to more pronounced clusters with better separation. To analyze this effect, we study the time evolution of the mean and variance of the clusters by plotting bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$ as time $n$ increases (cf. §2.4). Figure 3 shows the clustering into three clusters of a randomized first order to be picked from an initial randomized warehouse distribution, represented by probability density functions with covariance matrices $\Sigma_{1}(1),\Sigma_{1}(2),\Sigma_{1}(3)$, together with snapshots of the clusters for the $10$:th and $100$:th picking iteration of the same order, represented by probability density functions with covariance matrices $\Sigma_{n}(1),\Sigma_{n}(2),\Sigma_{n}(3)$ for $n=10$ and $n=100$, respectively. Each panel shows a superposition of the corresponding probability density functions. The figure shows that the clusters become more pronounced with better separation as time progresses.
Figure 3. Three snapshots of cluster distributions from experiment 1, illustrated by bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$, $\ell=1,2,3$. Each panel shows a superposition of the corresponding probability density functions. The snapshots are taken at time $n$ (i.e., in the language of §2.4 they show the clusters computed for the $n$:th order), with $n=1$ (left), $n=10$ (middle), $n=100$ (right). The orders are the same in every iteration. The figure clearly shows that as time progresses, the separation between cluster centers increases, while the bumps become more articulated due to a decreased variance. A movie showing the evolution of the clusters for $n=1,\ldots,200$ can be found here: https://play.hb.se/media/clusters_unperturbed/0_3r1qr1w4.
4.2. Experiment 2
To account for small differences in individual orders, we next consider the random example when $\omega=(o_{1},o_{2},\ldots)$ and the orders $o_{n}$ are independent identically distributed perturbations of the order $o$ used in the previous experiment. As described in §2.3 we then compute the warehouse state after picking the first $n$ orders via
$x_{n}=F_{\theta^{n-1}\omega}\circ\cdots\circ F_{\omega}(x_{0})=f(o_{n},\bullet)\circ\cdots\circ f(o_{1},\bullet)(x_{0}).$
We let each order $o_{n}$ be constructed as described under item 4 above, but with the first of the $10$ article types in each of the $20$ purchase orders randomly selected from the $890$ article types with uniform distribution. Clearly, two such purchase orders will differ by approximately $10\%$ in article types, and two consecutive picking orders consisting of $20$ such purchase orders that are mutually disjoint to begin with then also differ by approximately $10\%$ in article types, counting multiplicity. The resulting time evolution of clusters when picking and resorting such a sequence of orders $\omega=(o_{1},o_{2},\ldots)$ is described in Figure 4, and we can see that the resorting procedure still leads to clusters becoming more pronounced with better separation, even in the presence of small variations in orders. Note that since the initial clusters depend on both the initial warehouse state $x_{0}$ and the first order $o_{1}$, the initial clusters in this experiment will usually be different from the initial clusters in the previous experiment which were obtained from $x_{0}$ and the unperturbed order $o$.
Figure 4. Three snapshots of cluster distributions from experiment 2, illustrated by bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$, $\ell=1,2,3$. Each panel shows a superposition of the corresponding probability density functions. The snapshots are taken at time $n=1$ (left), $n=10$ (middle), $n=100$ (right). Two picking orders now differ by approximately $10\%$ on average. Although not as pronounced, we still clearly see an increased separation and decreased variance of the clusters as time progresses. A movie showing the evolution of the clusters for $n=1,\ldots,200$ can be found here: https://play.hb.se/media/clusters_static_perturbation/0_t9n5lbme.
4.3. Experiment 3
For comparison, we perform the same experiment but instead of always replacing the first of the $10$ article types in each purchase order with a random article type, we replace a randomly selected article type in each purchase order of the original order $o$ with a random article type. Two such purchase orders will then differ on average by around $19\%$ in article types, and two consecutive picking orders consisting of $20$ such purchase orders that are mutually disjoint to begin with also differ by approximately $19\%$ in article types on average, counting multiplicity. The result is described in Figure 5 and we can see that although still there, the clusters are now showing less of a tendency of becoming more pronounced with better separation.
Figure 5. Three snapshots of cluster distributions from experiment 3, illustrated by bivariate Gaussian probability density functions defined using each cluster’s mean and covariance matrix $\Sigma_{n}(\ell)$, $\ell=1,2,3$. Each panel shows a superposition of the corresponding probability density functions. The snapshots are taken at time $n=1$ (left), $n=10$ (middle), $n=100$ (right). Two picking orders now differ by approximately $19\%$ on average, and we only see a slight improvement in the separation and variance of the clusters as time progresses. A movie showing the evolution of the clusters for $n=1,\ldots,200$ can be found here: https://play.hb.se/media/clusters_random_perturbation/0_g0zx8ooj.
4.4. Quantitative comparison
As experiments 1, 2, and 3 show, the picking and resorting procedure leads in time to increased separation of the initial clusters, although this separation becomes less pronounced as the variability of the orders increases. To quantify this effect we measure the area of the triangle defined by the cluster center points, and see how this varies with time for the three different setups used in experiments 1, 2, and 3. In Figure 6 we show snapshots of the cluster centers for the clusters depicted in Figure 3 (taken at the same picking iterations $n=1,10,100$ as in Figure 3), and we can clearly see that the area is greatly increased with subsequent iterations. For experiment 1 the locations of the cluster centers vary fairly slowly, while for experiment 2 and especially 3 the locations vary more rapidly, and as a result the area fluctuates more for experiment 2 and 3. To increase signal strength we therefore average the calculations over 10 runs for each experiment type 1, 2, and 3, and compute the area as a function of time (number of iterations of the picking and resorting algorithm). The result is presented in Figure 7. For each experiment type, the average area increases from an initial average area of around $0.1$; for experiment 1 it has increased to around 6 after the first 20 iterations; for experiment 2 it has increased to around 3 after the first 20 iterations; for experiment 3 it has increased to around 1 after the first 20 iterations. In all cases the increase is immediately visible and of at least one order of magnitude in size.
Refer to caption222244446666888810101010222244446666888810101010
Refer to caption222244446666888810101010222244446666888810101010
Refer to caption222244446666888810101010222244446666888810101010
Figure 6. Cluster center points corresponding to the snapshots of the clusters in Figure 3 from experiment 1, with $x$ and $y$ coordinates on the axes. As in Figure 3, the snapshots are taken after $n=1$ (left), $n=10$ (middle), and $n=100$ (right) picking iterations. Movies showing how the location of the cluster centers change with every iteration $n=1,\ldots,200$ for the experiments depicted in Figures 3, 4, and 5 can be found here: https://play.hb.se/media/cluster_centers_unperturbed/0_5eassah3, https://play.hb.se/media/cluster_centers_static_perturbation/0_njd00ou2, and https://play.hb.se/media/cluster_centers_random_perturbation/0_lazlyssr.
Refer to caption010020000222244446666888810101010Time
Refer to caption010020000222244446666888810101010Time
Refer to caption010020000222244446666888810101010Time
Figure 7. Area of the triangle defined by the cluster center points as a function of the number of picking iterations (time) for experiment 1 (left), experiment 2 (middle), and experiment 3 (right). To increase signal strength, the area is averaged over 10 experiments of each type, shown in blue. The gray region indicates the average area plus minus one standard deviation. The average area for the initial warehouse state is around $0.1$, and in all three panels this increases substantially within the first 10 or 20 iterations as the clusters become more separated due to the picking and resorting procedure. However, the increase is visibly smaller for experiment 2 and even more so for experiment 3, indicating less pronounced separation of clusters as the noise level increases.
4.5. Route optimization
Here we analyze the effects that the resorting algorithm has on route optimization, and showcase the approximate method of calculating optimal routes that we describe in Section 3. We do so by analyzing an arbitrary route of a customer order, consisting of 10 products with varying quantities. The order is presented in Table 1. This particular order is assumed to be recurring (as in Experiment 1 above), and at the initial stage, iteration 1, the articles are randomly distributed throughout the entire domain of the warehouse node positions. Since the customer order is clustered, it will trigger a reorganization of clustered products, causing them to relocate closer to the cluster center. As stated earlier, the clusters tend to become more concentrated with reorganization, suggesting that the picking area within each cluster decreases. At iteration 300 we should therefore see a decrease in the total picking distance for each customer order.
Table 1. Products and their quantities. The position of each article in the warehouse is encoded through a value between 00 and $99$ to describe its $(x,y)$ coordinate in terms of a flattened $10\times 10$ grid starting from bottom left, together with a value between 00 and $9$ to describe its $z$ coordinate.
Product Quantity
a17_4 2
a36_6 2
a57_8 6
a9_7 6
a90_0 6
a77_8 7
a93_6 5
a38_1 4
a23_2 5
a24_4 4
When evaluating all possible route permutations through a graph with $n$ vertices, the total number of unique paths is given by $n!$, assuming each vertex is visited exactly once. For each such permutation, $(n-1)$ transitions must be evaluated to compute the total route cost, resulting in $n!(n-1)$ operations in total. In the CUDA C implementation, each route permutation may be assigned to a separate thread, and within each thread, an internal loop is used to process the transitions between vertices. If $V$ denotes the number of vertices (i.e., the number of stops per route), and $N=n!$ is the total number of permutations, then the computational structure involves operations proportional to $V\times N$. This scaling becomes problematic as $n$ increases, since $n!(n-1)$ grows factorially and dominates memory consumption. Empirically, we observed that values of $N$ larger than 2,903,040 lead to memory errors in the CUDA kernel, likely due to this exponential increase. As a result, processing large values of $n$ without segmenting the computation or distributing it across multiple kernel launches becomes infeasible due to the limitations in available GPU memory, see Figure 8.
Figure 8. Data from the Bellman Ford GPU version (RTX6000, 24GB VRAM) (left), and a comparison between $n!(n-1)$ and computational time (right).
Hence, since the number of stop positions in our experiment is $n=10$, we use the approximate method for calculating an optimal route described in Section 3. Initially, the number of stops is 10 while the number of passed nodes is 27, and the order is picked over a large portion of the warehouse area. By iteration 300, the clustered order regions have shrunk, and the number of passed nodes is now 15, determined using the approximate method. At iteration 300, the reorganization has resulted in many articles being located at the same stopping node, reducing the number of combinations to $6!$. This reduction enables a full combinatorial calculation of the optimal route. The result of the simulation is shown in Figure 9. As can be seen in this case, the reorganization yields an 80% increase in capacity. The approximate method does not produce the exact same result as the optimal method, with an optimal distance of 12. However, given the challenges with large $n!$, the approximate method allows for results close to the optimal solution in real-time when using GPUs.
Figure 9. Picking routes for the customer order from Table 1 at different iterations of the resorting algorithm. The stopping positions along the route are indicated by the larger nodes, and they are clustered into 3 color coded regions using the methodology from Section 3. At iteration 1 there are 10 stopping nodes, and the left panel shows the computed route. At iteration 300, the resorting algorithm has reduced the number of stopping positions to $6$, and the middle panel shows the computed approximately optimal route with length 15. For comparison, the right panel shows the optimal route at 300 iterations having length 12.
5. Discussion and conclusions
When we performed experiments of using unsupervised clustering to reduce computational complexity when determining routes as described at the end of Section 4, we often found that the computed route coincided with the optimal route after 300 iterations (as well as at earlier stages). However, since this is not guaranteed, we deliberately chose to present an experiment where this was not the case – the middle and right panels of Figure 9 are indeed different. In addition, when the separation between the different regions of stopping nodes (cf. the dashed rectangles in Figure 2) varies, some scenarios might call for a natural collapsing of two (perhaps somewhat overlapping) regions into one, thereby changing the clustering scheme. There is potential for analyzing the effects of such phenomena as well, but that is beyond the scope of this paper.
The implementation of the Bellman-Ford algorithm in CUDA C demonstrates that GPU-based parallelism is highly effective for solving routing problems, particularly when evaluating large numbers of path combinations. While the exponential growth of permutations presents a challenge, especially as the number of nodes increases, this limitation can be mitigated by dividing large routing problems into smaller sub-routes. By processing these segments independently, it is possible to maintain high computational efficiency without exceeding the GPU’s memory or thread capacity. This segmentation strategy ensures that GPU acceleration remains a scalable and practical solution even for complex graph traversal problems. Therefore, the combination of Bellman-Ford and CUDA C provides a flexible framework that balances algorithmic precision with hardware-aware execution.
The application of the mathematical models and numerical methods presented in this paper closely parallels contemporary advancements in industry practice. In particular, the dynamic clustering techniques described in our numerical experiments, where warehouse storage layouts adapt continuously based on changing order patterns, are exemplified in operations such as those in Amazon’s fulfillment centers. Amazon employs sophisticated robotic picking algorithms, dynamically optimizing the selection and clustering of inventory units to enhance operational efficiency and reduce picking times [1]. This real-world application aligns closely with our mathematical formulation of order clustering and its resulting impact on route optimization, underscoring the practical viability of our proposed approach. Our numerical experiments show that while random variability in orders reduces the clarity of clustering effects on the definition and refinement of picking regions, the overall impact remains significant. This suggests that our approach is robust enough for practical deployment in real-world environments characterized by inherent order variability.
Our exploration of GPU-accelerated route optimization algorithms, in particular the Bellman-Ford algorithm implemented through CUDA-based parallelization, also finds a clear reflection in industry. Alibaba’s implementation of vehicle routing algorithms, optimized through GPU computing techniques, exemplifies the practical benefits of this computational strategy, which has significantly enhanced logistics performance and reduced operational costs through improved route optimization [12]. The success of these implementations in a commercial context validates the computational approach proposed in our numerical experiments, demonstrating the scalability and effectiveness of GPU-based solutions for complex warehouse routing scenarios.
Appendix A WMS loop
Here we describe the algorithms of the WMS. The graph $G$ has all information about the position of each product, its current inventory balance, as well as empty positions that can be used for re-positioning a product so that it can be situated closer to a cluster center. (In the mathematical language of §2, the graph is thus a state $x\in X$.) The order graph consists of similar setup as the graph $G$, with the difference that this graph only has the order content (each article number) and the ordered amount of each product. (The order graph is thus an order $o$.) The WMS loop has one purpose, to enable orders that belong to a cluster to come closer to each other. This will in an ideal operating condition result in areas in the warehouse where only orders belonging to a specific cluster is situated, see Figure 1.
The Order Processing algorithm converts an order ID into a string representation and determines the position to test. It then updates the stock based on the order string and increments the pick count in the graph $G$.
Input: order_id, ordergraf
Output: order_string, test_pos
Function ProcessOrder($order\_id,ordergraf$):
       $order\_string,test\_pos\leftarrow order\_to\_string(order\_id,ordergraf)$
       UpdateStock($G,order\_string$)
       $increment\_pick\_count(G,order\_string)$
Algorithm 1 Order Processing
The Update Stock algorithm updates the stock in the graph $G$ based on the provided order string. It ensures that the inventory levels are adjusted according to the orders processed.
Input: G, order_string
Output: None
Function UpdateStock($G,order\_string$):
       $update\_stock\_from\_order(G,order\_string)$
Algorithm 2 Update Stock
The Check Stock algorithm verifies if there is sufficient stock in the graph $G$ to fulfill the next order string. It returns a Boolean value indicating whether the stock is sufficient.
Input: G, next_order_string
Output: Boolean
Function CheckStock($G,next\_order\_string$):
       return $is\_sufficient\_stock\_for\_order(G,next\_order\_string)$
Algorithm 3 Check Stock
The Check Stock and Move Articles algorithm first checks if there is sufficient stock for the next order. If the stock is insufficient, it identifies empty nodes and calculates the distances to determine the optimal movement of articles to balance the stock.
Input: G, ordergraf, order_id
Output: None
Function CheckStockAndMove($G,ordergraf,order\_id$):
       $next\_order\_string\leftarrow order\_to\_string(order\_id+1,ordergraf)$
      if CheckStock($G,next\_order\_string$) = False then
             $empty\_nodes\leftarrow\textnormal{{FindEmptyNodes(}}\textnormal{\emph{$G$}}\textnormal{{)}}$
             $specific\_point\leftarrow ordergraf.nodes[order\_id].get(^{\prime}cluster\_center^{\prime},^{\prime}Unknown^{\prime})$
             $blocked\_orders\leftarrow update\_stock\_from\_order2(G,next\_order\_string)$
             $distance\leftarrow\textnormal{{CalculateDistance(}}\textnormal{\emph{$empty\_nodes,specific\_point$}}\textnormal{{)}}$
             $min\_distance\leftarrow find\_min(distance)$
             $distance2\leftarrow\textnormal{{CalculateDistance(}}\textnormal{\emph{$blocked\_orders,specific\_point$}}\textnormal{{)}}$
            for $i\leftarrow 0$ to $length(distance2)-1$ do
                   $test\_pos\leftarrow distance2[i][1]>min\_distance[1]$
                  if $test\_pos$ then
                         MoveArticle($G,distance2[i][0],min\_distance[0],blocked\_orders[i][3]$)
                   end if
             end for
       end if
Algorithm 4 Check stock and Move Articles
The Main Order Processing Algorithm coordinates the overall process of handling orders. It iterates through a specified number of repetitions and processes each order by converting it to a string, updating the stock, and checking the stock to move articles as necessary.
Input: ordergraf, G, order_index, n_repetitioner
Output: None
Function Main:
       for $repetition\leftarrow 0$ to $n\_repetitioner-1$ do
             for $order\_id\leftarrow 1$ to $order\_index$ do
                   ProcessOrder($order\_id,ordergraf$)
                  if $order\_id<number\_of\_orders$ then
                         CheckStockAndMove($G,ordergraf,order\_id,order\_string$)
                   end if
             end for
       end for
Algorithm 5 Main Order Processing Algorithm